# Mamba-in-Mamba: Centralized Mamba-Cross-Scan in Tokenized Mamba Model for Hyperspectral Image Classification

### v1

@@ III-B Centralized Mamba-Cross-Scan (MCS)
The original HSI symbolized as $\mathbf{\mathbf{X}}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times C}$ can be defined as follows:
@@
where $\mathbf{\mathbf{x}}^{(i,j)}\in\mathbb{R}^{\mathcal{C}}$ signifies the spectral signature at the spatial position $(i,j)$. Consequently,
@@
where $32\leq i\leq\mathcal{H}-$ 3, $32\leq j\leq\mathcal{W}-$ 3, and $\mathbf{X}^{(i,\,j)}\in\mathbb{R}^{p\times p\times\mathcal{C}}$. The center pixel $\mathbf{X}^{(i,\,j)}[2,2]$ is $\mathbf{x}^{(i,\,j)}$.
Instead of the conventional Raster flattening used in other literatures, we adopt the Snake scan method (e.g., U-Turn scan), which has been proposed and validated in some research [36, 37, 38]. This method serves as the foundational scanning approach in our proposed MCS.
The Snake scan method can preserve the highest spatial continuity between pixels or partitioned windows when converting them into sequences. This approach minimizes the occurrence of jump connections when scanning reaches the boundaries of the image. Moreover, inspired by concept of multi-scanning strategy [38], which verifies that integrating multiple directional scanned sequences can augment sequential model training, we aim to extend it into this study.
Building on above background, we designed four types of centralized MCS that maximize the potential and lightweight when transforming the image into multi-directional sequences from multiple starting and ending points. The conceptual difference between Transformer’s self-attention mechanism and proposed MCS mechanism is shown in Fig. 4.
Specifically, as detailed in Fig. 5, each type of MCS begins from two opposing positions and converges at the center pixel, analogous to the half bi-directional processing found in RNNs but with centralized design.
For easy explanation below, we take the Eq. 9 as the example for better understanding below.
@@ III-B1 MCS Type-1
Figure 5: Four types of MCS with a $5\times 5$ patch. (a) Type-1, (b) Type-2, (c) Type-3, and (4) Type-4. They cover the all the possible continuous snake scanning for a patch.
Figure 6: The T-Mamba encoder. Take inputs of size $\mathbb{R}^{(5^{2})\times C_{2}}$ as an example. Initially, the projected input $\mathbf{s}$ is split into forward and backward sequences, each processed independently through a Conv1d and an S6 model. A Gaussian decay mask then statically adjusts the features of each sequence. Two processed feature sequences are then merged for feature aggregation. A downsample-driven Semantic Token Learner (STL) subsequently generates representative semantic tokens, which are correlated with $\mathbf{z}$ by an attention-driven Semantic Token Fuser (STF) to implement feature enhancement. Finally, a residual connection from the original input is used to enhance feature integration for output to the next stage.
@@ III-C Tokenized Mamba (T-Mamba)
After obtaining two sub-sequences from each type of MCS, we introduce the novel T-Mamba encoder, which processes them in a bi-directional manner inspired by the ViM [45]. An overview of the T-Mamba is shown in Fig. 6.
@@ III-C1 Whole Flowchart of T-Mamba
Taking Eq. 11 as an example, for a type-1 MCS, we start with an embedded sequence $\mathbf{S}_{1}^{(i,\,j)}\in\mathbb{R}^{5^{2}\times\mathcal{C}_{2}}$. This sequence is first normalized and then projected into two sequences as follows:
$\mathbf{s},\,\mathbf{z}\leftarrow{Linear}\Big{(}{Norm}\Big{(}\mathbf{S}_{1}^{(i,\,j)}\Big{)}\Big{)},$ (19)
where $\mathbf{s}$ and $\mathbf{z}$ are both of size $\mathbb{R}^{5^{2}\times\mathcal{C}_{3}}$.
Next, $\mathbf{s}$ is split into two directional sub-sequences, $\mathbf{s}_{f}$ and $\mathbf{s}_{b}$, where $f$ denotes the forward sub-sequence, and $b$ denotes the backward sub-sequence with each sub-sequence having dimensions $\mathbb{R}^{13\times\mathcal{C}_{3}}$, referring Eq. 12:
$\mathbf{s}_{f},\,\mathbf{s}_{b}\leftarrow{Split}\Big{(}\mathbf{s}\Big{)}.$ (20)
@@
Subsequently, these processed sub-sequences are input to the Mamba model as described in Sec. II:
@@ III-C2 Gaussian Decay Mask (GDM)
Instead of strictly utilizing binary values of 0 and 1 to determine which features to keep or discard, we propose the GDM as a soft mask that corresponds to the spectral and index distances in the sequence, centered on the pixel.
@@
$v_{t}=\exp\left(-\frac{1}{2}\left(\frac{\Big{|}\Big{|}\widetilde{\mathbf{s}_{f}}[t]-\widetilde{\mathbf{s}_{f}}[T]\Big{|}\Big{|}^{2}}{\sigma_{fea}}\right)^{2}\right)$ (33)
@@ III-C3 Cross Scan Merge
The ${Merge}()$ function is designed to seamlessly combine two sub-sequences by respecting their temporal dynamics, preparing them effectively for subsequent processing layers. We detail the computational steps below, as referenced in Eq. 24.
@@
where $\widehat{\mathbf{s}}$ resumes the original shape of the complete sequence, $\mathbb{R}^{p^{2}\times\mathcal{C}_{3}}$.
@@
Figure 7: The visualization process of STL.
@@ III-C4 Semantic Token Learner (STL)
Drawing inspiration from recent advancements [25, 54], we introduce a STL in this study. This component is designed to abstract and condense sequence input into a set of representative semantic tokens after processing through a sequential attention mechanism. This abstraction aids in downsampling and captures crucial high-level features for downstream tasks. Fig. 7 illustrates a sample of STL transformation process.
@@
Figure 8: The visualization process of STF.
@@ III-C5 Semantic Token Fuser (STF)
Next, the sequential attention $SeqAttn()$, mentioned in Eqs. 39a-41 is re-applied to $\overline{\mathbf{z}}$ to emphasize significant features:
@@
Figure 9: Three datasets, (a) Indian Pines (IP), (b) Pavia University (PU), and (c) Houston 2013 (HU2013) are used for experiments. The false-color images, fixed and disjoint training sets and testing sets are shown above.
TABLE I: Description of Land-Cover Types and Number of Train-Test Samples for Three Datasets
    Indian Pines (IP)     Houston 2013 (HU2013))     Pavia University (PU)
 No.     Class Name Color Train Test     Class Name Color Train Test     Class Name Color Train Test
1     Corn-notill 50 1,384     Grass-health 198 1,053     Asphalt 548 6,304
2     Corn-mintill 50 784     Grass-stress 190 1,064     Meadows 540 18,146
3     Corn 50 184     Grass-synthetic 192 505     Gravel 392 1,815
4     Grass-pasture 50 447     Tree 188 1,056     Trees 524 2,912
5     Grass-trees 50 697     Soils 186 1,056     Metal Sheets 265 1,113
6     Hay-windrowed 50 439     Waters 182 143     Bare Soil 532 4,572
7     Soybean-notill 50 918     Residential 196 1,072     Bitumen 375 981
8     Soybean-mintill 50 2,418     Commercials 191 1,053     Bricks 514 3,364
9     Soybean-clean 50 564     Roads 193 1,059     Shadows 231 795
10     Wheat 50 162     Highways 191 1,036
11     Woods 50 1,244     Railways 181 1,054
12     Buildings-Grass-Trees 50 330     Park-l-1 192 1,041
13     Stone-Steel-Towers 50 45     Park-l-2 184 285
14     Alfalfa 15 39     Tennis-ground 181 247
15     Grass-pasture-mowed 15 11     Running-track 187 473
16     Oats 15 5
 Total     695 9,671     2,832 12,197     3,921 40,002
Ratio     6.7% 93.3%     18.8% 81.2%     8.9% 91.1%
@@ III-D Weighted MCS Fusion (WMF)
TABLE II: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE INDIAN PINES DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
@@
III-E Multi-Scale Loss Design
IV Experimental results and analysis
In this section, we first describe three well-known HSI datasets [^5] http://lesun.weebly.com/hyperspectral-data-set.html. We then outline the implementation details and the methods used for comparison. Finally, we conduct extensive experiments to evaluate the performance of the proposed method both quantitatively and qualitatively.
IV-A Datasets Description
Three HSI datasets are illustrated in Fig. 9, along with their corresponding false-color images, training samples and testing samples. The numbers of training and testing samples for each class within these datasets are listed in Table I.
For fair comparison, all the training and testing samples are manually selected in a fixed and disjoint manner. Details can be referred to [18].
TABLE III: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE PAVIA UNIVERSITY DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
@@ IV-B Experimental Evaluation Indicators
TABLE IV: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE HOUSTON 2013 DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
@@
Figure 12: The classification maps with different methods for the HU2013 dataset. (a) vanilla ViT (OA=82.9548%). (b) SpeFormer (OA=86.4557%). (c) MAEST (OA=88.8415%) (d) HiT (OA=90.0303%). (e) SSFTT (OA=91.4241%). (f) 3DViT (OA=89.7024%). (g) HUSST (OA=90.1205%). (h) ViM (OA=87.5871 %). (i) MiM (OA=92.8917%).

### v3

@@ III-B Centralized Mamba-Cross-Scan (MCS)
The original HSI symbolized as $\mathbf{\mathbf{X}}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times\mathcal{C}_{0}}$ can be defined as follows:
@@
where $\mathbf{\mathbf{x}}^{(i,j)}\in\mathbb{R}^{\mathcal{C}_{0}}$ signifies the spectral signature at the spatial position $(i,j)$. Consequently,
@@
where $32\leq i\leq\mathcal{H}-$ 3, $32\leq j\leq\mathcal{W}-$ 3, and $\mathbf{X}^{(i,\,j)}\in\mathbb{R}^{p\times p\times\mathcal{C}_{0}}$. The center pixel $\mathbf{X}^{(i,\,j)}[2,2]$ is $\mathbf{x}^{(i,\,j)}$.
The existing centralized scanning method, SS2D, as utilized in [54], employs traditional Raster flattening across the spatial domain of images. This approach can lead to misalignment of image contents, particularly through jump connections that occur when the scan meets image boundaries. Furthermore, SS2D does not fully embrace the two-dimensional structure of the image, limiting its ability to analyze from multiple perspectives.
Consequently, we have opted for the Snake scan method (e.g., U-Turn scan), which has been endorsed in various studies [43, 44] and serves as the foundation for our proposed MCS methodology. The Snake scan method preserves spatial continuity between pixels or partitioned windows when transforming them into sequences, reducing the incidence of jump connections at image boundaries. This method’s effectiveness is further supported by the multi-scanning strategy concept [45], which demonstrates that using sequences scanned from multiple directions can enhance sequential model training.
Leveraging this background, we have developed four types of centralized MCS that optimize the transformation of images into sequences from various starting and ending points, ensuring both potential maximization and system efficiency. The fundamental distinction between the Transformer’s self-attention and our MCS mechanism is illustrated in Fig. 4.
As depicted in Fig. 5, each MCS type initiates from two opposing positions and converges at the center pixel. This method is reminiscent of the bi-directional processing typical of RNNs but incorporates a centralized design, providing a novel approach to image sequence processing.
Before feeding the $\mathbf{X}^{(i,\,j)}$ into MCS, we first deploy the depth-wise convolution (DWconv) [^3] https://github.com/rosinality/depthwise-conv-pytorch and point-wise convolution (PWconv) [^4] https://github.com/anshulpaigwar/Pointwise-Convolutional-Neural-Network on $\mathbf{X}^{(i,\,j)}$ for local feature aggregation and dynamic positional embedding.
$\mathbf{X}^{(i,\,j)}\leftarrow PWconv\Big{(}DWconv\Big{(}\mathbf{X}^{(i,\,j)}\Big{)}\Big{)}$ (10)
@@ III-B1 MCS Type-1
Figure 5: Four types of MCS with a $5\times 5$ patch. (a) Type-1, (b) Type-2, (c) Type-3, and (d) Type-4. They cover the all the possible continuous snake scanning for a patch.
Figure 6: The T-Mamba encoder. Take inputs of size $\mathbb{R}^{(5^{2})\times C_{2}}$ as an example. Initially, the input is split into forward and backward sequences $\mathbf{s}_{f}$ and $\mathbf{s}_{b}$ by MCS, each processed independently through a Conv1d and an S6 model. A Gaussian decay mask then statically adjusts the features of each sequence. Two processed feature sequences are then merged for feature aggregation. A downsample-driven Semantic Token Learner (STL) subsequently generates representative semantic tokens, which are correlated with $\mathbf{z}$ by an attention-driven Semantic Token Fuser (STF) to implement feature enhancement. Finally, a residual connection from the original input is used to enhance feature integration for output to the next stage.
@@ III-C Tokenized Mamba (T-Mamba)
After obtaining two sub-sequences from each type of MCS, we introduce the novel T-Mamba encoder, which processes them in a bi-directional manner. An overview of the T-Mamba is shown in Fig. 6.
@@ III-C1 Whole Flowchart of T-Mamba
Taking Eq. 12 as an example, for a type-1 MCS, we start with an embedded sequence $\mathbf{S}_{1}^{(i,\,j)}\in\mathbb{R}^{5^{2}\times\mathcal{C}_{2}}$. This sequence is first normalized and then projected into a sequence as follows:
$\mathbf{z}={Linear}\Big{(}{Norm}\Big{(}\mathbf{S}_{1}^{(i,\,j)}\Big{)}\Big{)},$ (20)
where $\mathbf{z}$ is of size $\mathbb{R}^{5^{2}\times\mathcal{C}_{3}}$.
Next, $\mathbf{S}_{1}^{(i,\,j)}$ is split into two directional sub-sequences, $\mathbf{s}_{f}$ and $\mathbf{s}_{b}$ by MCS, where $f$ denotes the forward sub-sequence, and $b$ denotes the backward sub-sequence with each sub-sequence having dimensions $\mathbb{R}^{13\times\mathcal{C}_{3}}$, referring Eq. 13:
$\mathbf{s}_{f},\,\mathbf{s}_{b}\leftarrow{MCS_{1}}\Big{(}\mathbf{S}_{1}^{(i,\,j)}\Big{)},$ (21)
where $MCS_{1}()$ denotes the split operation under type-1 MCS.
@@
Subsequently, these processed sub-sequences are input to the Mamba model as described in Section. II:
@@ III-C2 Gaussian Decay Mask (GDM)
To well distribute the influence of features from each step in the sequence, it is essential to apply nuanced weighting to each step’s features. Rather than relying on binary values of 0 and 1 to decide which features to retain or eliminate, we introduce the GDM as a soft mask. This GDM assigns weights based on spectral and index distances within the sequence, focusing around the central pixel.
@@
$v_{t}=\exp\left(-\frac{1}{2}\left(\frac{\Big{|}\Big{|}\widetilde{\mathbf{s}_{f}}[t]-\widetilde{\mathbf{s}_{f}}[T]\Big{|}\Big{|}^{2}}{\sigma_{fea}}\right)^{2}\right),$ (34)
where ||.$||.||^{2}$ represents the L2-distance (Euclidean distance).
@@
This approach may allow for a more refined control over feature selection, enhancing the model’s ability to capture and emphasize relevant spatial and spectral information.
@@ III-C3 Cross Scan Merge
To effectively merge features from two half-directional sequences and restore them to the original sequence length, we have designed the ${Merge}()$ function. This function is crafted to seamlessly integrate two sub-sequences, honoring their temporal dynamics and optimally preparing them for subsequent processing layers. The computational steps involved in this merging process are detailed below, as referenced in Eq. 25.
@@
where $\widehat{\mathbf{s}}$ resumes the original length of sequence, $\mathbb{R}^{p^{2}\times\mathcal{C}_{3}}$.
@@
Figure 7: The visualization process of STL. The merged sequence feature is denoted by $\widehat{\mathbf{s}}$. After applying sequential attention to $\widehat{\mathbf{s}}$, the resulting features are represented as $\widehat{\mathbf{m}}$. The learnable transformation matrices $\mathbf{U}_{1}$ and $\mathbf{U}_{2}$ are utilized to extract length-level and dimension-level semantic features, denoted by $\mathbf{A}$ and $\mathbf{V}$, respectively. Finally, the down-sampled semantic tokens $\mathbf{u}$ are derived through the inner product of $\mathbf{A}$ and $\mathbf{V}$.
@@ III-C4 Semantic Token Learner (STL)
Efficiently condensing coarse features into fine and dynamic semantic features within a sequence remains a question. Drawing inspiration from recent advancements [26, 65], we introduce the STL in our study. STL is tailored to abstract and condense sequence input into a set of representative semantic tokens, utilizing a sequential attention mechanism to enhance the focus and relevance of features. Fig. 7 depicts an example of the transformation process facilitated by the STL.
@@
This STL aids in downsampling and captures crucial high-level features for downstream tasks.
Figure 8: The visualization process of STF. It illustrates the semantic feature fusion of $\mathbf{z}$ and $\mathbf{u}$. To ensure compatibility in size with $\mathbf{u}$, $\mathbf{z}$ is first adaptively pooled into $\overline{\mathbf{z}}$. Subsequently, the pooled feature $\overline{\mathbf{z}}$ is processed through a learnable transformation matrix $\mathbf{Z}$ to produce $\widehat{\mathbf{z}}$, which captures the influence score. This allows the correlation between $\overline{\mathbf{z}}$ and $\mathbf{u}$ to be assessed through the inner product of $\widehat{\mathbf{z}}$ and the semantic token $\mathbf{u}$. Finally, to integrate the token-influenced data and the attention-modified data, $\overline{\mathbf{u}}$ and $\overline{\mathbf{m}}$ are summed to yield the final output $\widehat{\mathbf{u}}$.
@@ III-C5 Semantic Token Fuser (STF)
Next, the sequential attention $SeqAttn()$, mentioned in Eqs. 40a-42 is re-applied to $\overline{\mathbf{z}}$ to emphasize significant features:
@@
Figure 9: Four datasets, (a) Indian Pines (IP), (b) Pavia University (PU), (c) Houston 2013 (HU2013), and (d) WHU-Hi-HongHu (HongHu) are used for experiments. The false-color images, fixed and disjoint training sets and testing sets are shown above.
TABLE I: Description of Land-Cover Types and Number of Train-Test Samples for Four Datasets
    Indian Pines (IP)     Houston 2013 (HU2013))     Pavia University (PU)     WHU-Hi-HongHu (HongHu)
 No. Color     Class Name Train Test     Class Name Train Test     Class Name Train Test     Class Name Train Test
1     Corn-n. 50 1,384     Grass-h. 198 1,053     Asphalt 548 6,304     Red-r. 25 14,016
2     Corn-m. 50 784     Grass-s. 190 1,064     Meadow 540 18,146     Road 25 3,487
3     Corn 50 184     Grass-s. 192 505     Gravel 392 1,815     Bare-s. 25 21,796
4     Grass-p. 50 447     Tree 188 1,056     Tree 524 2,912     Cotton 25 163,260
5     Grass-t. 50 697     Soil 186 1,056     Metal-s. 265 1,113     Cotton-f. 25 6,193
6     Hay-w. 50 439     Water 182 143     Bare-s. 532 4,572     Rape 25 44,532
7     Soybean-n. 50 918     Residential 196 1,072     Bitumen 375 981     Chinese-c. 25 24,078
8     Soybean-m. 50 2,418     Commercial 191 1,053     Brick 514 3,364     Pakchoi 25 4,029
9     Soybean-c. 50 564     Road 193 1,059     Shadow 231 795     Cabbage 25 10,794
10     Wheat 50 162     Highway 191 1,036         Tuber-m. 25 12,369
11     Wood 50 1,244     Railway 181 1,054         Brassica-p. 25 10,990
12     Buildings-g.-t. 50 330     Park-l-1 192 1,041         Brassica-c. 25 8,929
13     Stone-s.-t. 50 45     Park-l-2 184 285         Small-b.-c. 25 22,482
14     Alfalfa 15 39     Tennis-g. 181 247         Lactuca-s. 25 7,331
15     Grass-p.-m. 15 11     Running-t. 187 473         Celtuce 25 977
16     Oat 15 5             Film-c.-l. 25 7,237
17                 Romaine-l. 25 2,985
18                 Carrot 25 3,192
19                 White-r. 25 8,687
20                 Garlic-s. 25 3,461
21                 Broad-b. 25 1,303
22                 Tree 25 4,015
Total     695 9,671     2,832 12,197     3,921 40,002     550 386,143
Ratio     6.7% 93.3%     18.8% 81.2%     8.9% 91.1%     0.14% 99.86%
TABLE II: The Hyperparameter settings and size of notations for Four datasets.
Definition     Notation     IP     HU2013     PU     HongHu
 Image Size:     ×$\mathcal{H}\times\mathcal{W}\times\mathcal{C}_{0}$     $145\times 145\times 200$     $349\times 1905\times 144$     $610\times 340\times 103$     $940\times 475\times 270$
Dimension after PCA:     $\mathcal{C}_{1}$     60     30     30     100
Initial Patch Size:     $p\times p$     $7\times 7$     $9\times 9$     $11\times 1$ 11     $9\times 9$
Feature Dimension Setting:     $\mathcal{C}_{2},\,\mathcal{C}_{3},\,\mathcal{C}_{4}$     64     64     32     128
Number of T-Mamba Layers:         4     4     2     3
Number of MiM Layers:         4     5     6     5
Learning Rate:         0.0005     0.0005     0.001     0.001
Feature Dropout:         0.1     0.1     0.2     0.1
@@ III-D Weighted MCS Fusion (WMF)
TABLE III: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE INDIAN PINES DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
@@
TABLE IV: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE PAVIA UNIVERSITY DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
@@
III-E Multi-Scale Loss Design
IV Experimental results and analysis
In this section, we first describe four well-known HSI datasets. We then outline the implementation details and the methods used for comparison. Finally, we conduct extensive experiments to evaluate the performance of the proposed method both quantitatively and qualitatively.
IV-A Datasets Description
Four HSI datasets are illustrated in Fig. 9, along with their corresponding false-color images, training samples and testing samples. The numbers of training and testing samples for each class within these datasets are listed in Table I.
The WHU-Hi-HongHu (HongHu) dataset [^8] http://rsidea.whu.edu.cn/resource_WHUHi_sharing.htm was acquired by the UAV platform on November 20, 2017, in Honghu City, Hubei Province, China. The experimental area is a complex agricultural scene with many classes of crops, and different cultivars of the same crop are also planted in the region. It consists of $940\times 475$ pixels and initially has 270 bands. This dataset has 22 different ground cover classes.
For fair comparison, all the training and testing samples are manually selected in a fixed and disjoint manner. Details can be referred to [19] and corresponding website.
@@ IV-B Experimental Evaluation Indicators
TABLE V: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE HOUSTON 2013 DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
@@
Figure 12: The region of interest (ROI) in classification maps with different methods for the HU2013 dataset. (a) vanilla ViT (OA=82.9548%). (b) SpeFormer (OA=86.4557%). (c) MAEST (OA=88.8415%) (d) HiT (OA=90.0303%). (e) SSFTT (OA=91.4241%). (f) 3DViT (OA=89.7024%). (g) HUSST (OA=90.1205%). (h) ViM (OA=87.5871 %). (i) MiM (OA=92.8917%).